OpenStudy (anonymous):

Solve. 3x^2 + 21x = –30 x = 7 and x = –5 x = –5 and x = –2 x = 6 and x = 5 x = –6 and x = –2

OpenStudy (akshay_budhkar):

-5,-2

OpenStudy (akshay_budhkar):

you want me to solve?

OpenStudy (anonymous):

first simplify by dividing by 3 x^2 + 7x + 10 = 0 (x+2)(x+5)=0 x = -5,-2

OpenStudy (akshay_budhkar):

jimmy did my job! thanks:)

OpenStudy (anonymous):

yw